Two, long, straight direction as shown. wires sepa
Answer #1

(a) Since the two wires carry current in same direction, they attract each other (the magnetic fields produced by these current will point in opposite directions. As a result the two wires attract each other)

(b) Magnetic field due to wire at a distance of 7cm is

                                                     B_1 = \frac{\mu _0i}{2\pi r} =\frac{4\pi *10^{-7}*3}{2\pi *0.07}=85.71*10^{-7}\:T

Magnetic field due to wire at a distance of 4cm is

                                                     B_2 = \frac{\mu _0i}{2\pi r} =\frac{4\pi *10^{-7}*3}{2\pi *0.04}=150*10^{-7}\:T

Total Magnetic field at Point P is = B1+B2 = 235.71*10^{-7} T

(b) using Pythogoras theorem,

          the first wire(first circled one) is at a distance of \sqrt{3^2+4^2} = 5 cm = 0.05 m

    Magnetic field due to wire at a distance of 5cm is

                                                     B_1 = \frac{\mu _0i}{2\pi r} =\frac{4\pi *10^{-7}*3}{2\pi *0.05}=120*10^{-7}\:T

Magnetic field due to wire at a distance of 4cm is

                                                     B_2 = \frac{\mu _0i}{2\pi r} =\frac{4\pi *10^{-7}*3}{2\pi *0.04}=150*10^{-7}\:T


B2 3, уст 9 3,Sine 2. 3 3cm 3 tan 53-130 Magneic Held 120 x10 coS 5330e Sin 53 130A